The following drugs were dissolved and prepared the 250 mL of stock solution.

Lidocaine: 12.6 mg, Procaineamide = 12.3 mg, Disopyramide = 12.7 mg, Quinidine = 12.2 mg

Standard 1: 25 mL of ~1mg/L of each drug diluted with mobile phase (0.5mL Stock solution + 24.5 mL mobile phase)

Standard 2: 25mL of ~2 mg/L of each drug diluted with mobile phase (1mL Stock solution +24 mL mobile phase)

Standard 3: 25mL of ~5 mg/L of each drug diluted with mobile phase (2.5 mLStock solution +22.5 mL mobile phase)

Standard 4: 25 mL of ~10mg/L of each drug diluted with mobile phase (5 mL Stock solution +20 mL mobile phase)

Standard 5: 25 mL of ~20 mg/L of each drug diluted with mobile phase (10 mL Stock solution + 15 mL mobile phase)

Question: Calculate the concentration of each drug in Stock A and Standards 1-5.
Answer #1

Lidocaine: 12.6 *4 = 50.4 mg/L

Procaineamide = 12.3 mg *4 = 49.2 mg/L

Disopyramide = 12.7 mg*4 = 50.8 mg/L

Quinidine = 12.2 mg*4 = 48.8 mg/L

------------------------------------------------------

Standerd A :

M1V1 = M2V2

where m1 and v1 is the concentration and volume of the stock. M2 and v2 is the concentration and volume of the diluted solution.

Lidocaine: (0.5 * 1mg/L) / 25 mL = 0.02 mg/L

Procaineamide = 0.02

Disopyramide = 0.02 mg/L

Quinidine = 0.02 mg/L

----------------------------------------------

Standard 2:

(1mL * 2mg/L )/25 mL = 0.08 mg/L

Standard 3 :

(2.5 mL *5mg/L)/25 mL = 0.5 mg/L

---------------------------------------------------

Standard 4;

(5 mL *10mg/L)/25mL = 2mg/L

----------------------------------------------------

Standard 5:

(10mL * 20mg/L )/25 mL = 8 mg/L